JONATHON77:

Find two consecutive integers whose sum is 113. Which of the following equations could be used to solve the problem?

shonyce15:

Let the two integers be x and x + 1, then x + x + 1 = 113 2x + 1 = 113

jhonyy9:

exactly these i ve wan writing too so what you need to do now again ?

jhonyy9:

solve this equation for x

jhonyy9:

@shonyce15 yes ? do you can solve it ?

shonyce15:

yes

jhonyy9:

ok than please do it for medal

shonyce15:

x=56

shonyce15:

or

shonyce15:

Step-by-step explanation: Let two consecutive integers are x and x+1 As per the statement: The sum of two consecutive integers is 113 ⇒x+x+1=113 Like terms are those which have same variable to the same power. Combine like terms in [1] we have; Therefore, the following equations could be used to solve the problem is,

jhonyy9:

yes 56 is one term and the second one what will be ?

jhonyy9:

56 + ? = 113

jhonyy9:

consecutive two integers

shonyce15:

2nd one would be 57

shonyce15:

56+57=113